OpenStudy (anonymous):

The slope of the line below is 3. Use the coordinates of the labeled point to find the point-slope equation of the line. (6,8) A. y - 6 = -3(x - 8) B. y - 8 = -3(x - 6) C. y - 8 = 3(x - 6) D. y - 6 = 3(x - 8)

OpenStudy (anonymous):

point slope form is \[y-y _{1}^{?}=m(x-x _{1}^{?})\] *ignore the Q marks

OpenStudy (anonymous):

What are the points on the line?

OpenStudy (anonymous):

(6,8)

OpenStudy (anonymous):

sub in the points on for y1 and x1, then sub 3 in for m. What'd ya get?

OpenStudy (anonymous):

hold on one sec

OpenStudy (anonymous):

ok sorry my dog wanted to go outside. so replace y with 6 and x with 8? or replace the 1's?

OpenStudy (anonymous):

ok this is something idk exactly how to do

OpenStudy (anonymous):

There are 2 diffrent xs and 2 diffrent ys in the point slope equation. They only want you to use one right now. What I was trying to say is. \[y _{1}=6\] \[x _{1}=8\] and \[m=3\] in \[y-y _{1}=m(x-x _{1})\] sub them and what do ya get?

OpenStudy (anonymous):

well 6-8 =-2

OpenStudy (anonymous):

Ohhhhh. wait it would be A.

OpenStudy (anonymous):

Sorry no all you have to do is substitute, not subtract. so y- ____=____(x-_____)

OpenStudy (anonymous):

y - 6 = -3(x - 8) So it wouldn't be that but replace y with 6 and x with 8?

OpenStudy (anonymous):

Yes